# LogicTree: Structured Proof Exploration for Coherent and Rigorous Logical Reasoning with Large Language Models

### v1

@@ 6 Conclusion
First, we evaluate our framework in the domain of logical reasoning, as it represents a distinct type of challenge in reasoning tasks that requires structured and extensive exploration. Our goal is to address this type of reasoning challenge, which often demands more deliberate reasoning. In future work, we plan to extend the framework to more complex domains such as theorem proving.
Second, our framework assumes that all premises (i.e., facts and rules) are explicitly provided in structured form. However, real-world scenarios often involve incomplete or unstructured knowledge. Future work may adapt our approach to scenarios where missing information must be retrieved from LLMs’ internal knowledge. Additionally, when facts and rules are not clearly separated, an extra pre-processing step with assistance from LLM may be required (Sun et al., 2024).
Third, our premise prioritization strategies rely on simple heuristics, which may not generalize well to more complex scenarios where inputs lack clear sentence structure. Developing rigorous strategies for proof planning and premise prioritization remains an important direction for future research.
@@
Our work adheres to the Code of Ethics. All utilized methods, models, and datasets are properly cited. The datasets used in our experiments are publicly available, and our research does not involve any private or sensitive information. We confirm that our use of datasets and LLMs aligns with their intended purposes and usage guidelines. For broader impacts, our framework contributes to the development of interpretable and automated reasoning systems. This line of work may extend to real-world applications that require rigorous, multi-step decision-making.
@@ Appendix A Computation for Semantic Overlap and Cumulative Connectivity
For each fact $f$, rule $r$, as well as the hypothesis $\mathcal{H}$, we obtain a triple:
$(f_{s},f_{p},f_{sp})$ from a fact,
$(r_{s},r_{p},r_{sp})$ from a rule,
$(\mathcal{H}_{s},\mathcal{H}_{p},\mathcal{H}_{sp})$ from the hypothesis,
where $s$, $p$, and $sp$ are extracted from the corresponding text using spaCy’s dependency parsing, with:
  • $s$ denoting the set of subjects,
  • $p$ denoting the set of predicates,
  • $sp$ denoting the set of subject-predicate pairs identified via parent-child relations in spaCy.
The semantic overlap $S(f,r)$ between a fact $f$ and a rule $r$ is defined as:
$S(f,r)=$ $0.25\cdot\mathbb{I}(f_{s}\cap r_{s}\neq\emptyset)$
+ $+\ 0.25\cdot\mathbb{I}(f_{p}\cap r_{p}\neq\emptyset)$
+ $+\ 0.5\cdot\mathbb{I}(f_{sp}\cap r_{sp}\neq\emptyset),$ (1)
and similarly, the semantic overlap $S(r,\mathcal{H})$ between a rule $r$ and hypothesis $\mathcal{H}$ is defined as:
$S(r,\mathcal{H})=$ $0.25\cdot\mathbb{I}(r_{s}\cap\mathcal{H}_{s}\neq\emptyset)$
+ $+\ 0.25\cdot\mathbb{I}(r_{p}\cap\mathcal{H}_{p}\neq\emptyset)$
+ $+\ 0.5\cdot\mathbb{I}(r_{sp}\cap\mathcal{H}_{sp}\neq\emptyset),$ (2)
@@
The cumulative connectivity $C(f,\mathcal{R})$ between a fact $f$ and the entire rule set $\mathcal{R}$ is defined as the sum of its semantic overlap with each rule in $\mathcal{R}$, i.e.,
$C(f,\mathcal{R})=\sum_{r\in\mathcal{R}}S(f,r).$
A higher cumulative connectivity value indicates that the fact is likely to initiate more reasoning branches through its relevant rules.
@@ C.1 Dataset
PrOntoQA-OOD  (Saparov et al., 2023) is a synthetic question-answering dataset using fictional characters. For evaluation, we use the most challenging 4-hop subset. We randomly sample 200 examples from the test set.
@@
Few-shot demonstrations for each LLM module are sampled from the training set of each dataset. An example of each dataset is shown in Appendix H.
@@ C.2 Models
All OpenAI models are accessed through OpenAI API [^2] https://platform.openai.com/docs/overview. Llama-3.3 70B is accessed through Together AI API [^3] https://www.together.ai/models/llama-3-3-70b. We set the temperature to 0.1 for all experiments. Our implementations leverage the Microsoft Guidance library [^4] https://github.com/guidance-ai/guidance.
@@ Appendix D Error Analysis on CoT, o1-mini, o3-mini, LogicTree
We manually conduct error analysis on CoT, o1-mini, o3-mini, and our framework using ProofWriter dataset. For CoT, we randomly sample 100 failed proofs, while for o1-mini, o3-mini, and our framework, we analyze all failed cases. We categorize the errors into three types: (1) insufficient exploration, (2) wrong derivation, and (3) hallucinated premise. The distribution of these error types for each method, along with illustrative examples, is shown in Figure 9. Our framework exhibits significantly fewer errors caused by insufficient exploration. In addition, it does not suffer from hallucinated premises, as access to the hypothesis is restricted to Verification Module only. This prevents the generation of unsupported premises that favor verifying the hypothesis during premise selection and inference.
@@
Appendix E Number of Reasoning Steps
@@
Table H shows the average number of reasoning steps on five logical reasoning datasets for different methods.
@@
The lower proof accuracy of LAMBADA (backward reasoning) compared to forward reasoning methods (ToT and LogicTree) can be explained by two main reasons: (1) LLMs demonstrate higher stepwise inference accuracy in forward reasoning as shown in Figure 3b; (2) the larger number of reasoning steps in backward reasoning increases the likelihood of making errors.
@@ Appendix G Using LLM Modules for Premise Prioritization
The prompts used for the LLM modules in fact and rule ranking are provided in Appendix H. We do not apply predefined criteria in these prompts, allowing us to assess LLM’s inherent ability on premise prioritization. For the average reasoning steps reported in Table 2, we do not include LLM queries related to fact and rule ranking. Even without this overhead, LLM-based premise prioritization results in more reasoning steps than our LLM-free heuristics. This reflects the limitation in LLM-based proof planning. Successfully guiding the reasoning process requires the model to already have an accurate understanding of how to reach the proof in advance, yet this ability is not evidenced by the limited performance of CoT (Table 3.4). Based on the results in Table 2, our simple LLM-free heuristics prove effective for strategic proof exploration.
Appendix H Dataset Example and Prompts for LLM Modules
@@
Figure 6: Schematic illustrations of baseline approaches to logical reasoning with LLMs: (a) Chain-of-Thought; (b) Selection-Inference; (c) Cumulative Reasoning; (d) Tree-of-Thought; (e) LAMBADA.
@@
3: Large language model $LLM$, NLP library spaCy
4:$s\leftarrow 0$triangleright Number of reasoning steps
5:$step\_limit\leftarrow L$
@@
9:         return $result$triangleright Verified
@@
12:$Fact\_Repo\leftarrow\mathcal{F}$
13:$Derivation\_HashMap\leftarrow\{\}$
14:$\mathcal{F}_{rank}\mathrel{\leftarrow}{\color[rgb]{0.0,0.0,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.7}\texttt{Fact\_Ranking}}(\texttt{spaCy},\mathcal{F})$
@@
16:     if $s>step\_limit$ then
@@
19:     $Tree\leftarrow[f_{i}]$triangleright Initiate the root of a tree
@@
21:         $f_{l}\leftarrow Tree.\texttt{pop()}$() $\triangleright$ LIFO, get the leaf fact
[…]
@@
25:         for each $r_{i}\in\mathcal{R}^{relv}_{rank}$ do $\triangleright$ Extend branches if $\mathcal{R}^{relv}\neq\emptyset$
@@
29:                  return $result$triangleright Verified
@@
32:                  $Tree.\texttt{append}(d_{i})$
@@
3: Large language model $LLM$
@@
[…]
@@
3: Large language model $LLM$
@@
8:     $\triangleright$ $result$ is either True or False
9:     return $result$
@@
Table 3: Complexity of premise search in CoT, ToT, LAMBADA, and our framework. An example involving conjunctive reasoning from ProofWriter is used for analysis. Note that although ToT’s complexity becomes linear in non-conjunctive cases, the reasoning type is not known beforehand. Therefore, ToT must retain its higher search complexity in the general case to accommodate complex scenarios.
@@
$m$ rules of the form If Dave is blue, then …,
$n$ rules of the form If Dave is cold, then …. The total number of paths to be explored is $m+n$.
@@
$m$ rules of the form If …, then Dave is blue,
$n$ rules of the form If …, then Dave is cold. The total number of paths to be explored is $m*n$.
Table 4: Analysis of three key factors that lead to more reasoning steps in backward reasoning compared to forward reasoning in iterative tree search for logical tasks: (1) Utilization of historical knowledge, (2) Evaluation of hypothesis and its negation, and (3) Branching complexity in reasoning paths. Together, these factors result in a higher average number of reasoning steps in backward reasoning, as illustrated in Figure 4 and Figure 8.
Figure 7: Proof accuracy of different methods across different levels of task difficulty (measured by depth) in ProofWriter. LogicTree consistently outperforms other methods at all depths, demonstrating superior robustness to problem difficulty.
@@
Table 5: Impact of the fact repository (fact repo) on proof accuracy within our framework.
@@
Figure 8: Proof accuracy vs. reasoning steps, averaged across five datasets for (a) GPT-4o-mini and (b) Llama-3.3 70B. The shaded area illustrates that our framework optimally scales inference-time computation to achieve higher proof accuracy.
Figure 9: Error analysis. (a)–(d) show the proportions of three error types for CoT (on GPT-4o), o1-mini, o3-mini, and LogicTree (on GPT-4o), respectively. (e) Examples of wrong inferences corresponding to the three error types.
@@
People who are round and red tend to be rough. If a person acts cold yet nice and green, they will be kind. If you meet someone with rough skin who is cold from being outside, you’ll notice they are nice. Every time you meet someone kind and nice, they’ll be green, too. Big people with red hair are cold because they cannot find coats that fit. It’s a certainty that any green, big and kind individual is going to be nice. A big round young person is often blue. Hypothesis:

### v2

@@ 6 Conclusion
First, we primarily evaluate our framework in the domain of logical reasoning, as it represents a distinct type of challenge in reasoning tasks that requires structured and extensive exploration. Our goal is to address this type of reasoning challenge, which often demands more deliberate reasoning. In future work, we plan to extend the framework to more complex domains such as theorem proving.
Second, our framework assumes that all premises (i.e., facts and rules) are explicitly provided. Future work could incorporate premise augmentation with plausible knowledge retrieved from LLM, rather than relying solely on the given premises. Additionally, when facts and rules are not clearly separated, an extra pre-processing step with assistance from LLM may be required (Sun et al., 2024). Also, our premise prioritization strategies rely on simple heuristics. Developing more advanced approaches for proof planning and premise prioritization remains an important direction for future research.
@@
Our work adheres to the Code of Ethics. All utilized methods, models, and datasets are properly cited. The datasets used in our experiments are publicly available, and our research does not involve any private or sensitive information. We confirm that our use of datasets and LLMs aligns with their intended purposes and usage guidelines. A potential risk of our framework lies in the misuse of its outputs in high-stakes domains without sufficient validation or expert review, as LLMs cannot always guarantee fully correct outputs. Nevertheless, when properly applied, our framework contributes to the development of interpretable and automated reasoning systems. Our work has the potential to extend to real-world applications that require rigorous, multi-step decision-making.
Acknowledgments
This work was supported in part by the Center for Co-Design of Cognitive Systems (CoCoSys), a Semiconductor Research Corporation (SRC) and DARPA-sponsored JUMP 2.0 center.
@@ Appendix A Computation for Semantic Overlap and Cumulative Connectivity
For each fact $f$, rule $r$, as well as the hypothesis $\mathcal{H}$, we extract a structured triple:
$(\mathcal{S}_{\text{Subj}}^{f},\mathcal{S}_{\text{Pred}}^{f},\mathcal{S}_{\text{SP}}^{f})$ from a fact,
$(\mathcal{S}_{\text{Subj}}^{r},\mathcal{S}_{\text{Pred}}^{r},\mathcal{S}_{\text{SP}}^{r})$ from a rule,
$(\mathcal{S}_{\text{Subj}}^{\mathcal{H}},\mathcal{S}_{\text{Pred}}^{\mathcal{H}},\mathcal{S}_{\text{SP}}^{\mathcal{H}})$ from the hypothesis.
In each triple:
  • $\mathcal{S}_{\text{Subj}}$ denotes the Set of Subjects,
  • $\mathcal{S}_{\text{Pred}}$ denotes the Set of Predicates,
  • $\mathcal{S}_{\text{SP}}$ denotes the Set of Subject-Predicate pairs identified via parent-child relations in spaCy.
These sets are extracted using spaCy’s dependency parser. We represent them as sets to account for the possibility of multiple subjects and predicates within a single fact, rule, or hypothesis.
The semantic overlap $Sem(f,r)$ between a fact $f$ and a rule $r$ is defined as:
$Sem(f,r)=\$$0.25\cdot\mathbb{I}(\mathcal{S}_{\text{Subj}}^{f}\cap\mathcal{S}_{\text{Subj}}^{r}\neq\emptyset)$
+ $+\ 0.25\cdot\mathbb{I}(\mathcal{S}_{\text{Pred}}^{f}\cap\mathcal{S}_{\text{Pred}}^{r}\neq\emptyset)$
+ $+\ 0.5\cdot\mathbb{I}(\mathcal{S}_{\text{SP}}^{f}\cap\mathcal{S}_{\text{SP}}^{r})\neq\emptyset),$ (1)
and similarly, the semantic overlap $Sem(r,\mathcal{H})$ between a rule $r$ and hypothesis $\mathcal{H}$ is defined as:
$Sem(r,\mathcal{H})=\$$0.25\cdot\mathbb{I}(\mathcal{S}_{\text{Subj}}^{r}\cap\mathcal{S}_{\text{Subj}}^{\mathcal{H}}\neq\emptyset)$
+ $+\ 0.25\cdot\mathbb{I}(\mathcal{S}_{\text{Pred}}^{r}\cap\mathcal{S}_{\text{Pred}}^{\mathcal{H}}\neq\emptyset)$
+ $+\ 0.5\cdot\mathbb{I}(\mathcal{S}_{\text{SP}}^{r}\cap\mathcal{S}_{\text{SP}}^{\mathcal{H}}\neq\emptyset),$ (2)
@@
The cumulative connectivity $\mathcal{C}(f,\mathcal{R})$ between a fact $f$ and the entire rule set $\mathcal{R}$ is defined as the sum of its semantic overlap with each rule in $\mathcal{R}$, i.e.,
$\mathcal{C}(f,\mathcal{R})=\sum_{r\in\mathcal{R}}Sem(f,r).$ (3)
A higher cumulative connectivity value indicates that the fact $f$ is likely to initiate more reasoning branches through its relevant rules.
@@ C.1 Dataset
We evaluate on five English-language logical reasoning datasets, as detailed below:
@@
PrOntoQA-OOD  (Saparov et al., 2023) is a synthetic question-answering dataset using fictional names. For evaluation, we use the most challenging 4-hop subset. We randomly sample 200 examples from the test set.
@@
Few-shot demonstrations for each LLM module are sampled from the training set of each dataset. An example of each dataset is shown in Appendix I.
@@ C.2 Models
All OpenAI models are accessed through OpenAI API [^3] https://platform.openai.com/docs/overview. Llama-3.3 70B is accessed through Together AI API [^4] https://www.together.ai/models/llama-3-3-70b. We set the temperature to 0.1 for all experiments to encourage more deterministic generation. All results are obtained from a single run. We utilize the Microsoft Guidance library in our implementation [^5] https://github.com/guidance-ai/guidance.
The version of spaCy model used in our framework is en_core_web_lg-3.8.0 (382 MB).
@@ Appendix D Error Analysis on CoT, o1-mini, o3-mini, LogicTree
We manually conduct error analysis on CoT, o1-mini, o3-mini, and our framework using ProofWriter dataset. For CoT, we randomly sample 100 failed proofs, while for o1-mini, o3-mini, and our framework, we analyze all failed cases. We categorize the errors into three types: (1) insufficient exploration, (2) wrong derivation, and (3) hallucinated premise. The proportion of these error types for each method, along with illustrative examples, is shown in Figure 10. Our framework exhibits significantly fewer errors caused by insufficient exploration. In addition, our framework does not suffer from hallucinated premises, as access to the hypothesis is restricted to Verification Module only. This prevents the generation of unsupported premises that favor verifying the hypothesis during premise selection and inference.
@@
Appendix E Number of Reasoning Steps, Generated Tokens, and Inference Time
The following elaborates on how we measure the number of reasoning steps for each approach.
@@ Appendix E Number of Reasoning Steps, Generated Tokens, and Inference Time
Table 6 shows the average number of reasoning steps across five logical reasoning datasets for different methods. Table 7 presents the average number of generated tokens and inference time for those different methods. The number of generated tokens is obtained using completion_tokens from the completion response.
@@
The lower proof accuracy of LAMBADA (backward reasoning) compared to forward reasoning methods (ToT and LogicTree) can be explained by two main reasons: (1) LLMs demonstrate higher stepwise inference accuracy in forward reasoning as shown in Figure 3b (§ 5.1); (2) the larger number of reasoning steps in backward reasoning increases the likelihood of making errors.
@@ Appendix G Using LLM Modules for Premise Prioritization
The prompts used for the LLM modules in fact and rule ranking are provided in Appendix I. We do not apply predefined criteria in these prompts, allowing us to assess LLM’s inherent ability on premise prioritization. For the average reasoning steps (LLM-based prioritize) reported in Table 2, we do not include LLM queries related to fact and rule ranking. Even without this overhead, LLM-based premise prioritization results in more reasoning steps than our LLM-free heuristics. This reflects the limitation in LLM-based proof planning. Successfully guiding the reasoning process requires the model to already have an accurate understanding of how to reach the proof in advance, yet this ability is not evidenced by the limited performance of CoT (Table 1). Based on the results in Table 2, our simple LLM-free heuristics prove effective for strategic proof exploration.
Appendix H Extension to Mathematical Reasoning
LogicTree is primarily designed to strengthen the logical reasoning capabilities of LLMs. Beyond its original focus, it can be readily adapted to other types of reasoning tasks that start from a given set of information. Our framework systematically combines relevant information for derivation, leverages derived information, and facilitates structured problem solving. In Figure 11, we illustrate how LogicTree performs mathematical reasoning on an example from GSM8K (Cobbe et al., 2021). We further evaluate and compare the accuracy of CoT, ToT, and LogicTree on 300 randomly sampled examples from GSM8K test set and 300 randomly sampled examples from MathQA (Amini et al., 2019) test set. All methods are evaluated on Qwen2.5-7B (Yang et al., 2024), with LogicTree demonstrating superior performance.
Appendix I Dataset Example and Prompt for LLM Modules
@@
Figure 6: Schematic illustrations of baseline approaches for logical reasoning with LLMs: (a) Chain-of-Thought; (b) Selection-Inference; (c) Cumulative Reasoning; (d) Tree-of-Thought; (e) LAMBADA.
@@
3: Large language model LLMLLM, NLP library spaCy
[…]
@@
9:   return resultresult $\triangleright$ Verified
@@
[…]
@@
16:  if s>step_limits>step\_limit then
@@
19:  $Tree\leftarrow[f_{i}]$ $\triangleright$ Initiate the root of a tree
@@
21:   $f_{l}\leftarrow Tree.\texttt{pop()}$ $\triangleright$ LIFO, get the leaf fact
22:   $\mathcal{R}^{relv}\leftarrow{\color[rgb]{0.8,0.4,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0.0}\texttt{Forward\_Selection}}($
@@
25:   for each $r_{i}\in\mathcal{R}^{relv}_{rank}$ do
@@
29:      return resultresult $\triangleright$ Verified
@@
32:      $Tree.\texttt{append}(d_{i})$ $\triangleright$ Extend branches
@@
3: Large language model LLMLLM
@@
12:  $f_{s}\leftarrow{\color[rgb]{0.8,0.4,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0.0}\texttt{Backward\_Selection}}($
@@
3: Large language model LLMLLM
@@
8:  $\triangleright$ resultresult is either True or False
9:  return resultresult
@@
Table 3: Complexity of premise search in CoT, ToT, LAMBADA, and LogicTree. An example involving conjunctive reasoning from ProofWriter is used for analysis. Note that although ToT’s complexity becomes linear in non-conjunctive cases, the reasoning type is not known beforehand. Therefore, ToT must retain its higher search complexity in the general case to accommodate complex scenarios.
@@
$m$ rules of the form: If Dave is blue, then …
$n$ rules of the form: If Dave is cold, then … The total number of paths to be explored is $m+n$.
@@
$m$ rules of the form: If …, then Dave is blue.
$n$ rules of the form: If …, then Dave is cold. The total number of paths to be explored is mnm*n.
Table 4: Analysis of three key factors that lead to more reasoning steps in backward reasoning compared to forward reasoning in iterative tree search for logical reasoning tasks: (1) Utilization of historical knowledge, (2) Evaluation of hypothesis and its negation, and (3) Branching complexity in reasoning paths. Together, these factors result in a higher average number of reasoning steps in backward reasoning, as illustrated in Figure 4 and Figure 8.
@@
Table 5: Impact of the fact repository (fact repo) on proof accuracy across five datasets within our framework.
Figure 7: Proof accuracy of different methods across different levels of task difficulty (measured by depth) in ProofWriter on GPT-4o. LogicTree consistently outperforms other methods at all depths, demonstrating superior robustness to problem difficulty.
Figure 8: Proof accuracy versus reasoning steps, averaged across five datasets for (a) GPT-4o-mini and (b) Llama-3.3 70B. The shaded area illustrates that our framework optimally scales inference-time computation to achieve higher proof accuracy.
Figure 9: Proof accuracy versus token budget for different methods (SI, ToT, LogicTree) on ProofWriter and RobustLR datasets, evaluated with GPT-4o.
@@
Method Dataset
LogicNLI ParaRules PrOntoQA-OOD ProofWriter RobustLR
Token # Time (min) Token # Time (min) Token # Time (min) Token # Time (min) Token # Time (min)
CoT 145.4 0.11 105.5 0.08 125.9 0.09 103.2 0.08 171.9 0.12
SI 681.7 0.80 688.3 0.79 233.2 0.28 490.5 0.68 502.1 0.60
CR 1231.6 1.43 871.6 1.04 490.4 0.62 816.7 0.94 800.7 0.97
LAMBADA 1181.0 1.36 758.6 0.97 566.3 0.67 550.5 0.74 921.4 1.07
ToT 1896.7 2.13 258.4 0.30 568.2 0.70 709.3 0.86 355.6 0.42
o1-mini 3253.3 1.11 464.5 0.19 943.0 0.39 953.1 0.42 534.8 0.24
o3-mini 2412.3 1.02 589.2 0.27 683.0 0.34 849.1 0.36 481.9 0.21
LogicTree 3627.0 4.07 440.6 0.54 1082.0 1.47 738.4 0.89 493.9 0.61
Table 7: Average number of generated tokens and inference time (in minutes) for different methods across five logical reasoning datasets. CoT, SI, CR, LAMBADA, ToT, and LogicTree results are based on GPT-4o.
Figure 10: Error analysis. (a–d) show the proportions of three error types for CoT (on GPT-4o), o1-mini, o3-mini, and LogicTree (on GPT-4o), respectively. (e) Examples of wrong proof corresponding to the three error types.
Figure 11: LogicTree for mathematical reasoning. Left: Case study illustrating how LogicTree solves a math problem from GSM8K. Right: Accuracy comparison among CoT, ToT and LogicTree on GSM8K and MathQA.
@@
People who are round and red tend to be rough. If a person acts cold yet nice and green, they will be kind. If you meet someone with rough skin who is cold from being outside, you’$l$ notice they are nice. Every time you meet someone kind and nice, they’$l$ be green, too. Big people with red hair are cold because they cannot find coats that fit. It’s a certainty that any green, big and kind individual is going to be nice. A big round young person is often blue. Hypothesis: